Company purchases 8000 units of a product each year at a unit cost of $10.00. The order cost is $30.00 per order, and the holding cost per unit per year is $3.00.

What are the economic order quantity, the total annual cost, the number of orders to place in one year, and the reorder point when the lead time is 2 weeks?

The supplier is now offering a special discount and has temporarily reduced his unit price from $10.00 to $9.00.

What amount should be purchased to take advantage of the discount?
Answer #1

A

EOQ = (2*annual demand*ordering cost/holding cost per unit per year)^.5

EOQ = (2*8000*30/3)^.5

EOQ = 400 units

B.

Total annual cost (without product cost) = (8000/400)*30 + (400/2)*3 = $1200

Total annual cost (with product cost) = 8000*10 + (8000/400)*30 + (400/2)*3 = $81200

C.

No. of orders to be placed in one year = 8000/400 = 20 orders

D

Weekly demand = (8000/52)

Reorder point = (8000/52)*2 = 307.7 or 308 units

Pl. repost other unanswered questions for their proper answers!